OpenStudy (anonymous):

In your own words, explain how a trinomial of the form 2x2 + 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping. Use complete sentences.

OpenStudy (cwrw238):

first multiply 2 by 15 = 30 now look for factors of 30 and look to see if you can get the 13 : 30 = 2 * 15 30 = 3 * 10 3 + 10 = 13 so we write: 2x^2 + 10x + 3x + 15

OpenStudy (cwrw238):

so we gruop the first 2 and last two terms to give 2x( x + 5) + 3(x + 5) = (2x + 3)(x + 5)

OpenStudy (cwrw238):

hope this helps

OpenStudy (anonymous):

thanks